OpenStudy (nickthurmm1):

If the average of six numbers is 18 and the average of three of the numbers is 15, then what is the sum of the remaining three numbers?

OpenStudy (nickthurmm1):

a. 21 c. 53 b. 45 d. 63

OpenStudy (photon336):

\[\frac{ \sum~numbers }{ amount } = average \] \[\frac{ 108 }{ 6 } = 18\] if we've got 6 numbers and the average is 18 then the sum must be 108 because 108/6 = 18. if the average of three of the numbers is 15 then \[\frac{ 45 }{ 3} =15 \] so to find out we do this: 108-43 = ?

OpenStudy (photon336):

@nickthurmm1